OpenStudy (anonymous):

how to find the exact value of tan^-1(tan 7π/4)

OpenStudy (jojokiw3):

Well you need to find tan(7pi/4) first, which you can use the unit circle for. https://www.mathsisfun.com/geometry/images/circle-unit-radians.gif Remember also that \[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\]

OpenStudy (aishalicious1):

Step 1)Find ur phone .... Step 2) Google it Step 3 )If that doesnt work whip out ur reliable calc push the right buttons and pray u pass My method genius isnt it? Its a miracle I pass this crap.